OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) = x2 g(x) = x2 + 20x + 100 In which direction and by how many units should f(x) be shifted to obtain g(x)?

OpenStudy (anonymous):

@abb0t Help please?

OpenStudy (anonymous):

@ankit042 Helo please?

OpenStudy (jdoe0001):

do you know what "perfect square trinomial is?

OpenStudy (jdoe0001):

\(\bf f(x) = x^2\qquad \qquad g(x) = x^2 + 20x + 100\) notice g(x) closely

OpenStudy (anonymous):

Yes, f(x) is part of g(x) @jdoe0001

OpenStudy (jdoe0001):

right... do you know what "perfect square trinomial is?

OpenStudy (anonymous):

I do it has three parts, one sqared, one x and a regular number @jdoe0001

OpenStudy (jdoe0001):

well... yes... well, g(x) is a "perfect square trinomial" so you should be able to make it into a squared binomial like \(\bf (a + b)^2\)

OpenStudy (jdoe0001):

if you were to do so, what would g(x) look like?

OpenStudy (anonymous):

Oh so I just factor? @jdoe0001

OpenStudy (jdoe0001):

yes, just factor it, once you factor, you'll notice the "shifts"

OpenStudy (anonymous):

Oh wow okay! Could you help me with another one? @jdoe0001

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x2 + 6x + 6 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 2 to x = 5 represent? @jdoe0001

OpenStudy (ankit042):

dx/dt represents the speed in this case

OpenStudy (jdoe0001):

... right

OpenStudy (anonymous):

I dont understand what im supposed to do @jdoe0001 @ankit042

OpenStudy (jdoe0001):

the average "rate of change" will involve the derivative which is what ankit042 said, the dx/dt

OpenStudy (ankit042):

find f(2) and f(5) now rate of change = (f(5)-f(2))/(5-2)